# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v2

@@ 3.3 Relative predictive power reveals different control timescales across modalities
Figure 4: Predictive power across input modalities and contexts. A Model predictions for lateral foot placement using CoM-relevant kinematics during treadmill walking, shown across increasing input time windows. B Explained variance ($R^{2}$, median and interquartile range) as a function of gait phase for modality-based (orange) and baseline (blue) prediction. The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. C Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. D Relative predictive power of CoM-relevant kinematics for treadmill walking. E Relative predictive power of CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
3.4 The swing foot dynamics predict future foot placement mid-swing
@@
4 Discussion
In this study, we present a data-driven framework to characterize real-world motor behaviors by how much and when different input modalities predict future actions. By employing deep learning models that incorporate the temporal history of state estimates, we outperform simpler models in non-stationary tasks for predicting future foot placements. We compare the predictive power of various hypothesized input modalities (e.g., center-of-mass kinematics, full-body kinematics, and gaze fixations) against a swing foot prediction baseline. This baseline comparison reduced the impact of inter-trial variability, allowing us to isolate and interpret the context- and modality-dependent effects more clearly. We quantify the contribution of each modality to foot placement through their distinct prediction timescales. We discover a tradeoff between long- and short-timescale predictions across contexts; in more uncertain contexts like rough terrain, we find a reduced reliance on long timescale predictions. Finally, rather than assuming that foot placement is initiated at the very start of the swing cycle, our method helps identify the swing phase at which the foot begins to carry significant predictive information about the foot placement location. Together, these findings highlight multi-timescale control strategies in real world locomotion, revealing how they are shaped by the task context and the input modality. Our framework can be extended to analyze other input modalities and motor actions to characterize real world motor behavior.
@@ 4 Discussion
Our work highlights how complex real world movement data can be combined with data-driven approaches to obtain insight into the control strategies at play. We benchmark various deep learning models, discovering that nonlinear action prediction models outperform traditional linear models by capturing the nonlinear input history-dependence at play in real world contexts. Our framework provides a comprehensive and generalizable approach to compare the influence of different input modalities and contexts on the control strategies chosen. By identifying specific phases when there is more predictive ability, it provides new insight into the temporal structure of motor planning and control in high-dimensional tasks. Overall, we present a behavioral data-driven approach that future work can build on to understand context- and modality-dependent control strategies in everyday movements. Our approach can be extended to analyze large-scale data across different motor behaviors [74, 75] and species [76, 77]. Integrating the context- and modality-dependent timescales identified by our approach into physics-based models could yield more human-like simulation and clarify the underlying neuromechanics [78, 79, 80]. These data-driven models can also enable more human-like assistive robots that generalize better across environmental contexts or more directly incorporate gaze data to inform control.
@@ Acknowledgments
To ensure a fair comparison, we independently tune the hyperparameters for each model and input modality across all contexts using nested cross-validation [81, 82]. See section 2.3 for details. For hyperparameter optimization, we use grid search when the search space is manageable, and random search [83] when the hyperparameter space exceeds 100 configurations. Below, we outline the hyperparameter search space for each model. See section 2.2.1 for details on model architecture.
@@
Section 2.3 describes how the performances of the models are evaluated.
Figure 6: Predictive power across models. A Predictive power during treadmill walking. The vertical line indicates the critical phase when the relative predictive power is maximized, thereby only gait phases before the gray-shaded area are considered. The model with the average highest score prior to the critical phase was deemed optimal, and its predictive power is shown by the interpolated curve (LOWESS). See section 2.3 for more details on the evaluation metric. The normalized model score is summarized in Table 1. B Predictive power during treadmill running.

### v3

@@ Results
Figure 4: Predictive power across input modalities and contexts. A Model predictions for lateral foot placement using CoM-relevant kinematics during treadmill walking, shown across increasing input time windows. B Explained variance ($R^{2}$, median and interquartile range) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) prediction. The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. C Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. D Relative predictive power of CoM-relevant kinematics for treadmill walking. E Relative predictive power of CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
The swing foot dynamics predict future foot placement mid-swing
@@
Discussion
In this study, we present a data-driven framework to characterize real-world motor behaviors by how much and when different input modalities predict future actions. We incorporate trial embedding information into the framework in an architecture-dependent manner, which enables the models to separate trial-specific explained variance from the variance explained by the input modality. By employing deep learning models that incorporate the temporal history of state estimates, we outperform simpler models in non-stationary tasks for predicting future foot placements. We compare the predictive power of various hypothesized input modalities (e.g., center-of-mass kinematics, full-body kinematics, and gaze fixations) against an autoregressive baseline. This baseline comparison reduced the impact of inter-trial variability, allowing us to isolate and interpret the context- and modality-dependent effects more clearly. We quantify the contribution of each modality to foot placement through their distinct prediction timescales. We discover a tradeoff between long- and short-timescale predictions across contexts; in more uncertain contexts like rough terrain, we find a reduced reliance on long timescale predictions. Finally, rather than assuming that foot placement is initiated at the very start of the swing cycle, our method helps identify the swing phase at which the foot begins to carry significant predictive information about the foot placement location. Together, these findings highlight multi-timescale control strategies in real world locomotion, revealing how they are shaped by the task context and the input modality. Our framework can be extended to analyze other input modalities and motor actions to characterize real world motor behavior.
@@ Discussion
Our work highlights how complex real world movement data can be combined with data-driven approaches to obtain insight into the control strategies at play. We put forth a deep learning framework for action prediction, finding that nonlinear models outperform traditional linear models by capturing the input history-dependence at play in real world contexts. Our framework provides a comprehensive and generalizable approach to compare the influence of different input modalities and contexts on the control strategies chosen. By identifying specific phases when there is more predictive ability, it provides new insight into the temporal structure of motor planning and control in high-dimensional tasks. Overall, we present a behavioral data-driven approach that future work can build on to understand context- and modality-dependent control strategies in everyday movements. Our approach can be extended to analyze large-scale data across different motor behaviors [74, 75] and species [76, 77]. Integrating the context- and modality-dependent timescales identified by our approach into physics-based models could yield more human-like simulation and clarify the underlying neuromechanics [78, 79, 80]. These data-driven models can also enable more human-like assistive robots that generalize better across environmental contexts or more directly incorporate gaze data to inform control.
@@ Acknowledgments
To ensure a fair comparison, we independently tune the hyperparameters for each model and input modality across all contexts using nested cross-validation [81, 82]. For hyperparameter optimization, we use grid search when the search space is manageable, and random search [83] when the hyperparameter space exceeds 100 configurations. Below, we outline the hyperparameter search space for each model. See the Methods section for details on model architecture.
@@
The Methods section describes how the performances of the models are evaluated.
Figure 6: Predictive power across models. A Predictive power during treadmill walking. The vertical line indicates the critical phase when the relative predictive power is maximized, thereby only gait phases before the gray-shaded area are considered. The model with the average highest score prior to the critical phase was deemed optimal, and its predictive power is shown by the interpolated curve (LOWESS). See the Methods section for more details on the evaluation metric. The normalized model score is summarized in Table 1. B Predictive power during treadmill running.